TEX SYMBOLS Large Operators

∑ \sum  ⋂ \bigcap  ⨀ \bigodot
∏ \prod  ⋃ \bigcup  ⨂ \bigotimes
∐ \coprod  ⨆ \bigsqcup  ⨁ \bigoplus
∫ \int  ⋁ \bigvee  ⨄ \biguplus
∮ \oint  ⋀ \bigwedge

Binary Relations

± \pm  ∩ \cap  ∨ \vee, \lor
∓ \mp  ∪ \cup  ∧ \wedge, \land
\ \setminus  ⊎ \uplus  ⊕ \oplus
· \cdot  ⊓ \sqcap  ⊖ \ominus
× \times  ⊔ \sqcup  ⊗ \otimes
∗ \ast  ◁ \triangleleft  ⊘ \oslash
⋆ \star  ▷ \triangleright  ⊙ \odot
⋄ \diamond  ≀ \wr  † \dagger
∘ \circ  ◯ \bigcirc  ‡ \ddagger
• \bullet  ÷ \div  ⨿ \amalg
△ \bigtriangleup  ▽ \bigtriangledown

Relations

≤ \leq, \le  ≥ \geq, \ge  ∼ \sim
≺ \prec  ≻ \succ  ≃ \simeq
⪯ \preceq  ⪰ \succeq  ≍ \asymp
≪ \ll  ≫ \gg  ≈ \approx
⊂ \subset  ⊃ \supset  ≅ \cong
⊆ \subseteq  ⊇ \supseteq  ∝ \propto
⊑ \sqsubseteq  ≠ \neq  ⊧ \models
∈ \in  ∋ \ni, \owns  ≐ \doteq
⊢ \vdash  ⊣ \dashv  ⊥ \perp
≡ \equiv  ⊣ \dashv  ▷◁ \bowtie

Remark \not negates most relations.

Arrows

← \leftarrow, \gets  ⟵ \longleftarrow  ↑ \uparrow
⇐ \Leftarrow  ⟸ \Longleftarrow  ⇑ \Uparrow
→ \rightarrow, \to  ⟶ \longrightarrow  ↓ \downarrow
⇒ \Rightarrow  ⟹ \Longrightarrow ⇓ \Downarrow
↔ \leftrightarrow  ⟷ \longleftrightarrow ↕ \updownarrow
⇔ \Leftrightarrow  ⟺ \Longleftrightarrow,\iff ⇕ \Updownarrow
↦ \mapsto  ⟼ \longmapsto  ↗ \nearrow
↪ \hookleftarrow  ↩ \hookrightarrow  ↘ \searrow
↼ \leftharpoonup  ⇀ \rightharpoonup  ↙ \swarro
↽ \leftharpoondown ⇁ \rightharpoondown ↖ \nwarrow

⇀ ↽ \rightleftharpoons Openings and Closings

[ \lbrack  { \lbrace  ⌊ \lfloor  ⟨ \langle  ⌈ \lceil
] \rbrack  } \rbrace  ⌋ \rfloor  ⟩ \rangle  ⌉ \rceil

Miscellaneous Symbols

ℵ \aleph  h̄ \hbar  ı \imath
ȷ \jmath  ℓ \ell  ℘ \wp
ℜ \Re  ℑ \Im  ∂ \partial
∞ \infty  ′ \prime  ∅ \emptyset
∇ \nabla  √ \surd  ⊤ \top
⊥ \bot  | \|, \vert  ∥ \∥, \Vert
\angle  △ \triangle  \ \backslash
∀ \forall  ∃ \exists  ¬ \neg, \lnot
♭ \flat  ♮ \natural  ♯ \sharp
♣ \clubsuit  ♢ \diamondsuit  ♡ \heartsuit
♠ \spadesuit  § \S  ¶ \P

Math Accents

â \hat a  à \grave a  ā \bar a
ǎ \check a  ȧ \dot a  a⃗ \vec a
ã \tilde a  ä \ddot a  âb \widehatab
á \acute a  ă \breve a  ãb \widetildeab